# Bubble in bubbles

### v1

@@ III Dynamic of double-layered bubbles
where $p=\Delta V_{12}~\text{or}~\Delta V_{23}$ denotes the pressure difference across the bubble wall.
To better understand the relationship between wall velocity and bubble radius, we simplify the equation of motion for the bubble wall. The Lorentz factor $\gamma$ characterizes the velocity of the bubble wall in relativistic terms, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$, and it accounts for the relativistic effects as the bubble expands. For an expanding bubble, the initial radius must exceed its critical radius. This condition can be solved analytically, with the initial condition $\gamma(R_{0})=1$, as derived in Ref. [30]. The equation is given by
$\gamma=\frac{pr}{3\sigma}+\frac{R_{0}^{2}}{r^{2}}-\frac{pR_{0}^{3}}{3\sigma r^{2}}\approx\frac{2r}{3R_{0}}+\frac{R_{0}^{2}}{3r^{2}},$ (11)
where in the last step, we assume that the initial radius is only slightly larger than the critical one, i.e., $R_{0}\approx R_{c}$. As shown in Fig. 5, the parameter $\alpha$ directly affects the potential differences $\Delta V_{12}$ and $\Delta V_{23}$, leading to distinct evolutionary outcomes. This can be understood from Eq. (11): the acceleration of a bubble wall is determined by the pressure difference across it. A larger potential difference drives the outermost wall to move faster, while the inner wall propagates more slowly. In contrast, a negative $\alpha$ can cause the inner wall to overtake the outer wall, resulting in a collision and subsequent merging of the two walls. Conversely, if the outer wall expands faster than the inner wall, the vacuum bubble can stably maintain a double-wall structure. Such a configuration may enhance or amplify wall-related physical phenomena, including gravitational wave production and baryogenesis.
@@
Having discussed how the relative velocities of the inner and outer walls govern the stability of a single double-layered bubble, we now consider the scenario in which two such bubbles nucleate close enough to interact. Assuming that the bubbles are aligned along the $z$-axis, the system exhibits approximate cylindrical symmetry, which allows us to simplify the analysis of their subsequent evolution. In this setup, the dynamics of the scalar field during the collision can be described by the Klein–Gordon equation in cylindrical coordinates:
@@
The walls of these classically formed elliptical bubbles subsequently collide with the inner walls of the original bubbles, leading to the formation of so-called trapping regions [31, 32], where trapping at the false vacuum occurs after the collisions. The formation of these trapping regions depends sensitively on the wall thickness and typically occurs in collisions involving thin-walled bubbles. Later, interactions between trapping regions and inner regions at $t/D\simeq 4.73$ further convert trapped $\phi_{2}$ regions into the $\phi_{3}$ vacuum. As time progresses, all vacua will eventually settle into the $\phi_{3}$ state by the scalar radiation. From our numerical results, we find that the energy–momentum tensor associated with the dynamics of double-layered bubbles exhibits significant differences from that of standard single-wall bubble collisions. Consequently, these differences are expected to lead to distinct features in the power spectrum of gravitational wave energy density.
@@ IV Conclusion and discussion
In this work, we have investigated flyover vacuum decay in a multi-vacuum scenario. We find that, for certain choices of initial parameters, this process can lead to the formation of double-layered vacuum bubbles. In contrast, double-layered bubbles generated purely by quantum tunneling are generally unstable and quickly decay into single-wall vacuum bubbles.
Our study employs a semiclassical approach originally introduced by Braden et al. [15] , who used the stochastic method to study vacuum decay numerically in 1+1 dimensions and demonstrated quantitative agreement with standard tunneling calculations. This raises the question of whether this semiclassical approach merely provides an alternative approximation for the decay rate or actually describes a distinct channel of vacuum decay. From our analysis of bubble nucleation in multi-vacuum potentials, the existence of stable double-layered bubbles suggests that flyover transitions indeed provide an independent decay channel, different from conventional quantum tunneling.
Furthermore, the dynamics of double-layered bubbles differ significantly from single-wall bubbles. The expansion and collisions of these bubbles modify the scalar field evolution, which in turn can drive the evolution of coupled fluids, potentially leading to distinct cosmological signatures such as gravitational waves or baryogenesis. Additionally, the presence of two types of nucleated bubbles introduces multiple possible decay pathways, giving rise to gravitational wave power spectra that depend on a variety of parameters controlling the bubble dynamics. Overall, our results indicate that semiclassical flyover transitions enrich the phenomenology of vacuum decay in multi-vacuum potentials, opening new possibilities for observational consequences in the early Universe.
@@ acknowledgements
This work is supported in part by the National Key Research and Development Program of China No. 2020YFC2201501, in part by the National Natural Science Foundation of China No. 12475067 and No. 12235019.

### v2

@@ III Dynamics of double-layered bubbles
where $p$ is the pressure difference across the bubble wall, and $\sigma$ represents the wall tension. This thin-wall description is expected to become increasingly accurate once the wall becomes relativistic, since the wall thickness is Lorentz-contracted in the lab frame.
To better understand the relationship between the wall velocity and the bubble radius, we simplify Eq. (10). The Lorentz factor $\gamma$ characterizes the wall velocity, defined as $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. It can be solved analytically with the initial condition that the wall is at rest. The solution is given by [15]
$\gamma=\frac{pr}{3\sigma}+\frac{R_{c}^{2}}{r^{2}}-\frac{pR_{c}^{3}}{3\sigma r^{2}}\approx\frac{2r}{3R_{c}}+\frac{R_{c}^{2}}{3r^{2}},$ (11)
where we have assumed that the initial radius is only slightly larger than the critical radius $R_{c}$. As shown in Fig. 5, the parameter $\alpha$ directly affects the potential differences $\Delta V_{12}$ and $\Delta V_{23}$, leading to distinct evolutionary outcomes. This can be understood from Eq. (11): the velocity of a bubble wall is determined by the pressure difference across it. A larger potential difference drives the outer wall to move faster, while the inner wall propagates more slowly. In contrast, a negative $\alpha$ can cause the inner wall to overtake the outer wall, resulting in a collision and subsequent merging of the two walls. Conversely, if the outer wall expands faster than the inner wall, the vacuum bubble can stably maintain a double-wall structure.
@@
Furthermore, when two double-layered bubbles collide, they can each destabilize the other’s inner and outer walls. Assuming that the bubbles are aligned along the $z$-axis, the system exhibits approximate cylindrical symmetry, which allows us to simplify the analysis of their subsequent evolution. In this setup, the dynamics of the scalar field during the collision can be described by the Klein–Gordon equation in cylindrical coordinates
@@
The walls of these classically formed elliptical bubbles subsequently collide with the inner walls of the original bubbles, leading to the formation of so-called trapping regions [22, 18], where trapping at the false vacuum occurs after the collisions. The formation of these trapping regions depends sensitively on the wall thickness and typically occurs in collisions involving thin-walled bubbles. Later, interactions between trapping regions and inner regions at $t/D\simeq 4.73$ further convert trapped $\phi_{2}$ regions into the $\phi_{3}$ vacuum. As time progresses, all vacua will eventually settle into the $\phi_{3}$ state by the scalar radiation. From our numerical results, we find that the energy–momentum tensor associated with the dynamics of double-layered bubbles can differ significantly from that of standard single-wall bubble collisions. While we do not compute gravitational-wave spectra in this work, existing studies indicate that the characteristic frequency scale is controlled by macroscopic length scales (e.g. the mean bubble separation / typical bubble size at collision), whereas the ultraviolet (high-frequency) fall-off can be sensitive to microscopic wall-structure effects such as the wall thickness; therefore the additional scales introduced by the double-layer dynamics may imprint qualitatively new spectral structure (e.g. a shoulder/secondary feature or a modified UV slope), which we leave for dedicated GW simulations.
@@ IV Conclusion and discussion
In this work, we have studied flyover vacuum decay in a multi-vacuum system, demonstrating that it can produce double-layered vacuum bubbles. We have examined their stability, showing that the interplay between inner and outer wall velocities determines whether the double-layered structure is preserved or merged. Additionally, collisions between such bubbles can further disrupt the configuration. Our findings shed light on the dynamics of non-standard vacuum decay and may have implications for early-universe phase transitions and associated observable signatures.
The flyover vacuum decay differs fundamentally from conventional quantum tunneling. In tunneling-induced vacuum decay, bubbles typically nucleate sequentially in time, with their spatial locations distributed randomly. By contrast, flyover vacuum decay proceeds by overcoming successive potential barriers, leading to overlapping bubble centers and the emergence of double-layered structures. Our analysis of bubble nucleation in multi-vacuum potentials shows that the existence of stable double-layered bubbles indicates flyover transitions constitute an independent decay channel, distinct from standard quantum tunneling.
It was theoretically shown that stepwise tunneling processes can generate nested bubbles [27, 11], where new bubbles nucleate inside older ones as the transition proceeds. In such models, flyover vacuum decay can also occur, enhancing the number of bubbles formed at each stage. However, since the multi-vacuum structure arises from two or more fields, the flyover process cannot traverse successive barriers continuously, but it can still overcome individual barriers, thereby producing configurations that resemble nested bubbles. Unfortunately, although both mechanisms may coexist during the phase transition, their dynamical processes are indistinguishable.
Furthermore, the dynamics of double-layered bubbles differ significantly from single-wall bubbles. The expansion and collisions of these bubbles modify the scalar field evolution, which in turn can drive the evolution of coupled fluids, potentially leading to distinct cosmological signatures such as gravitational waves or baryogenesis. Additionally, the presence of two types of nucleated bubbles introduces multiple possible decay pathways, giving rise to gravitational wave power spectra that depend on a variety of parameters controlling the bubble dynamics. Overall, our results indicate that flyover transitions enrich the phenomenology of vacuum decay in multi-vacuum potentials, opening new possibilities for observational consequences in the early Universe.
@@ acknowledgements
This work is supported in part by the National Natural Science Foundation of China No. 12235019 and No. 12475067.